OpenStudy (anonymous):

Still having trouble with this: Find a constant c between 1 and 9 such that the average value of the function f(x) on the interval [1,9] is equal to f(c). I've been told to split the integral, but what next?

OpenStudy (anonymous):

\[f(x)=6x^2-4x+2\]

OpenStudy (amistre64):

find the value of (f(9)-f(1))/(9-1) i believe to get the average value this tells us the slope between the points (1,f(1)) and (9,f(9))

OpenStudy (amistre64):

once we know that average value of the fucntion we can define that for the value of the function and solve for an appropriate x - which will be c

OpenStudy (anonymous):

So, I don't need to use integrals? This is an integral lesson...

OpenStudy (amistre64):

.... well, it is quite possible that my interpretation of the question is off :)

OpenStudy (anonymous):

Should we use \[1 \div (b-a)*\int\limits_{1}^{9}f(x)\]?

OpenStudy (anonymous):

That's the average value equation...

OpenStudy (anonymous):

oops..1=a and 9=b

OpenStudy (amistre64):

f(1) = 4 ; f(9) = 9( 6(9) -4) +2 = 9(50) + 2 = 452 452-4 448 ------ = ---- = 56 9-1 8 6x^2 -4x +2 = 56 ; when x= (1 +- sqrt(82))/3

OpenStudy (amistre64):

les see if that pans out :)

OpenStudy (anonymous):

Well, neither of those are choices. it's multiple choice.

OpenStudy (amistre64):

\[\frac{1}{8}\int\limits_{1}^{9}6x^2-4x+2\ dx\] \[\frac{1}{8}(2.9^3-2.9^2+2.9-2.1^3+2.1^2-2.1)\] \[\frac{1}{8}(2.9^3-162+18 -2)=164\]

OpenStudy (anonymous):

That's an answer. So, where does the f(c) come in...can you explain that?

OpenStudy (amistre64):

I spose you want a value of c that makes f(x)=164

OpenStudy (anonymous):

Oh! That makes a lot of sense. So, 6c^2-4c+2=164?

OpenStudy (anonymous):

You are the best! Thank you!

OpenStudy (amistre64):

yw

OpenStudy (amistre64):

i still think the question needs to be reworded tho :)